# Systematic Reward Gap Optimization for Mitigating VLM Hallucinations

### v3

@@ 1 Introduction
To address these limitations, we propose Topic-level Preference Rewriting (TPR), a novel VLM hallucination mitigation paradigm designed for systematically optimizing reward gap configuration during data curation, as illustrated in Figure 1 (a). To provide precise control over semantic details in the response and enable fine-grained contrast between $y_{w}$ and $y_{l}$, TPR operates at the topic-level. Specifically, TPR first decomposes responses into semantic units and clusters them into distinct topics. It then performs intra-topic self-resampling using the model itself to avoid introducing external biases that may occur with rewriting-based methods. Preference pairs $(y_{w},y_{l})$ are subsequently constructed by selectively replacing original semantic units with these alternatives from the same topic. This core mechanism allows for precise, fine-grained adjustment of semantic details between $y_{w}$ and $y_{l}$, guided by various chosen strategies. For instance, a greedy strategy within TPR can establish highly discriminative reward gaps by constructing $y_{w}$ and $y_{l}$ using the highest- or lowest-scored alternatives from their respective topics.
Moreover, this flexible control over fine-grained differences uniquely enables the formulation and investigation of sophisticated strategies for achieving an optimal reward gap configuration. In this work, we exemplify this capability with a simple yet effective curriculum learning strategy that progressively adjusts the difficulty of hallucinations included in the $y_{l}$. By implementing such hard negative mining, the model can be trained to more effectively counteract subtle and challenging hallucinations, showcasing the effectiveness of TPR. Crucially, this effectiveness is coupled with superior data efficiency, stemming directly from TPR’s ability to curate high-quality preference pairs. This enhanced data efficiency, clearly depicted in Figure 1 (b), underscores how TPR’s systematic optimization of reward gap configuration leads to more performant and cost-effective VLM alignment.
@@ 3.1 Preliminary: Preference Data and Alignment
Mitigating visual hallucinations in VLMs often involves aligning a policy model $\pi_{\theta}$ with carefully curated preference data, denoted as $\mathcal{D}=\{(I,x,y_{w},y_{l})\}$. Typically, for a given image $I$ and prompt $x$, these methods employ a base reference model $\pi_{\text{ref}}$ to generate candidate responses. Human experts or proxy AI labelers $\pi_{\text{label}}$ then evaluate these responses or rewrite them to form preference pairs $(y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$. The target policy model $\pi_{\theta}$, often initialized from $\pi_{\text{ref}}$, is subsequently fine-tuned on $\mathcal{D}$ using Direct Preference Optimization (DPO) [21]. DPO aligns $\pi_{\theta}$ with an reward function $r(y;x)$ that reflects the preference probability $p(y_{w}\succ y_{l}|x)$ inherent in the curated data. As highlighted by recent studies [20, 23, 24], the effectiveness of such alignment critically depends on the true reward gaps, formally defined as the difference $r(y_{w};x)-r(y_{l};x)$ within each preference pair. The strategic configuration of reward gaps across the curated preference data provides substantial training signals that not only reinforce desired behaviors but also accurately expose the model’s genuine and harder-to-discern deficiencies. Therefore, to sculpt effective reward gaps, two key principles should guide data curation: (1) The process must enable fine-grained control over data characteristics in $y_{w}$ and $y_{l}$ to deliberately shape reward gaps. (2) The preference pairs should accurately reflect desired behaviors while exposing model’s intrinsic failure modes.
@@
Guided by these principles, we propose Topic-level Preference Rewriting (TPR), a novel paradigm designed to systematically optimize reward gap configuration by establishing fine-grained control over topic-level semantic details within the response. TPR implements this through two core steps: topic-level alternatives generation (Section 3.2) and selective topic replacement (Section 3.3). Building upon this precise control, TPR enables the active exploration and strategic shaping of reward gap configuration, exemplified by the curriculum learning strategy detailed in Section 3.4.
@@ 3.2 Topic-level Alternatives Generation
VLM responses comprise various semantic topics, encompassing diverse objects and attributes, intricate spatial relationships, or subtle contextual implications. To offer flexible control over fine-grained details within these responses, TPR operates precisely at the topic level. Such a topic-centric approach is made viable by findings like those in RLAIF-V [18], which suggest topics within a response often exhibit weak correlations, permitting their relatively independent manipulation. As illustrated in Figure 2, TPR introduces a topic-level alternative generation approach designed to provide a rich set of candidates for the subsequent selective replacement of individual semantic topics.
@@
Topic Clustering. To facilitate subsequent selective replacement of semantic units with alternatives under the same topic, we group semantically related units from different candidate responses into unique topic clusters, based on both textual consistency and visual correlation. Textual consistency between units $u_{m,n}$ and $u_{p,q}$ is assessed by directly querying the reference model $\pi_{\text{ref}}$, e.g., “Are $u_{m,n}$ and $u_{p,q}$ describing the same topic?”. For visual correlation, we utilize features from the VLM’s visual encoder (e.g., CLIP [45]) to verify whether the units refer to similar regions within the input image $I$. This visual check is essential for disambiguating textually similar units that describe visually distinct entities. Units $u_{m,n}$ and $u_{p,q}$ are considered the same topic only if they satisfy both textual consistency and visual correlation criteria. We then apply a greedy connectivity algorithm [46] to perform topic clustering, yielding topic cluster $\{c\}$, each containing textually and visually related units. More details are provided in Appendix.
@@ 3.3 Selective Topic Rewriting
Figure 3: Constructing Preference Pairs by Selectively Replacement. Starting with a pool of scored topic alternatives, preference pairs are constructed through selectively replacing units within the template. The specific alternatives are chosen based on strategies like greedy or curriculum learning one, to deliberately control over the resulting reward gap. In-context rewriting is then employed to seamlessly integrate these chosen alternatives, ensuring fluent preference pairs for subsequent model alignment.
@@ 3.4 Exploring Optimal Reward Gap Configuration
Leveraging this control, we propose a simple yet effective curriculum learning strategy, analogous to hard negative mining in other domains, to optimize the reward gap configuration for hallucination reduction. This strategy involves structuring the preference data curation and DPO training in stages. Specifically, during an initial “Warm-Up” stage, we adopt the greedy strategy described above, constructing $(y_{w},y_{l})$ pairs with maximal score divergence between corresponding semantic units. This provide a strong initial learning signal, encouraging faster convergence by clearly differentiating between highly discriminative content. In a subsequent “Hard-Mining” stage, we gradually increase the difficulty of the learning objective by constructing $y_{l}$ using incorrect alternative units that have progressively higher scores, i.e., they are less obviously wrong and closer to the decision boundary. Introducing these “hard negative” examples in $y_{l}$ challenges the model to make finer distinctions during DPO training. By learning to differentiate preferred responses from subtly flawed ones, the model can more effectively refine its ability to detect and avoid nuanced hallucinations, ultimately improving robustness. This curriculum strategy systematically shape the overall reward gap configuration over time, starting with broad distinctions and moving towards finer-grained ones ultimately. Further details are provided in Appendix.
@@ 4.1 Experimental Setup
Models. In line with previous studies, we use LLaVA-1.5-7B [7] as both the reference model $\pi_{\text{ref}}$ for generating preference data, and as the policy model $\pi_{\theta}$ that is subsequently fine-tuned. The labeler model $\pi_{\text{label}}$, used for intra-topic ranking described in Section 3.3, is LLaVA-NeXT-34B [4], a choice consistent with recent approaches [18, 28]. In our experiments, we evaluate two variants of proposed TPR: (1) TPR utilizes the greedy replacement strategy for data curation. (2) TPR-CL (Curriculum Learning) utilizes a curriculum learning strategy for data curation, detailed in Section 3.4. We compare TPR against a range of counterparts, including ranking-based methods [16, 18, 19, 20, 25] and rewriting-based methods [13, 27, 28] that utilize either human experts [13] or external models [27, 28] for rewriting.
@@
[…]
@@
Data Source. Following common practice [13, 18], we curate our preference data based on 7 publicly available dataset sources: VQA v2 [47], MSCOCO [48], ShareGPT-4V [49], TextVQA [50], MovieNet [51], OKVQA [52] and Google Landmark v2 [53]. We ultimately generate a total of 20,000 preference data instances used for alignment. For the TPR-CL variant, 12,000 instances (60%) are constructed during “Warm-Up” stage, and the remaining 8,000 (40%) are constructed during “Hard-Mining” stage. More details are provided in Appendix.
@@ 4.2 Main Results
Results in Table 1 demonstrate TPR significantly mitigates visual hallucinations in VLMs across several benchmarks. The proposed TPR, in both its evaluated variants, achieves leading performance in mitigating visual hallucinations when applied to LLaVA-1.5-7B. Notably, our TPR-CL variant establishes new state-of-the-art results on several hallucination benchmarks, reducing hallucinations by $\sim$93on Object-HalBench and $\sim$41on MMHal-Bench. This substantial improvements can be consistently observed across other challenging hallucination benchmarks such as AMBER and RefoMB, underscoring the efficacy of our proposed design. Moreover, models fine-tuned using TPR-curated data not only maintain but sometimes enhance the base model’s performance on general capability benchmarks. This outcome indicates that TPR can effectively suppresses visual hallucinations without compromising general VLM capabilities.
@@
[…]
@@
[…]
@@ 4.3 Ablation Studies
To investigate how TPR optimizes the reward gap configuration by enabling deliberate control over underlying data characteristics within preference pairs, we conduct a series of ablation studies. As presented in Table 2, our investigations focus on addressing the following questions: (1) How do the core components of TPR influence the quality and data characteristics of the resulting preference pairs, which directly shape the individual reward gaps? (2) How does the selective replacement mechanism in TPR impact final performance of the learned policy, when compared against other preference curation approaches, such as ranking-based or rewriting-based methods? (3) How to construct a high-quality rejected response $y_{l}$ in preference pair? Is it more effective to employ external “black-box” rewriters common in rewriting-based methods or to utilize model’s internally resampled candidates, as TPR does? (4) From a broader perspective, what is the data efficiency of preference data curated by TPR, particularly when comparing its performance with that of other methods across varying data volumes? For computational efficiency, these ablation studies are conducted using a subset of 8,000 curated preference data instances. More ablations are provided in the Appendix, including using different model architectures, using the reference model itself as labeler, computational overhead, analysis on mitigated hallucinations and qualitative case studies.
Topic Alternatives Generation. We conduct ablations on different approaches for generating topic alternatives within the TPR. The results in Table 2 reveal that employing multiple initial candidate responses for decomposition and subsequently performing intra-topic self-resampling proves highly beneficial (exp 3a vs.3e/f). This approach facilitates better coverage of edge cases, such as rarely mentioned attributes, and ultimately yields a more diverse and comprehensive set of alternatives. Such diversity is crucial for sculpting more sophisticated reward gap configuration during data curation. To quantify these aspects, Figure 4 (b) provides a direct comparison, evaluating the underlying data characteristics like informativeness and trustworthiness of these generated alternatives by different approaches, based on GPT-4V reviews.
Preference Curation Approaches. We conduct ablations to analyze the impact of specific design choices within the preference curation process, as detailed in Table 2. Deviations from our base TPR variant (exp 3a), such as employing random selection for replacement (exp 3b), applying selective replacement to only one side of preference pair (exp 3c, or omitting in-context rewriting (exp 3d), all degrade final model performance. These findings highlight that the purposeful, symmetric replacement for both preferred and rejected responses, coupled with in-context rewriting, is crucial for crafting reward gaps that provide strong and effective learning signals for DPO.
Moreover, TPR’s core mechanism of selective topic replacement demonstrably outperforms other preference curation strategies, such as ranking-based or rewriting-based methods (exp 3a vs.3g/h/i/j). These performance advantages are further substantiated by quality assessments present in Figure 4 (a). In this evaluation, responses generated by different curation strategies are compared against strong “ground truth” (GT) responses from GPT-4V. We calculate one-vs.-one win-rates to quantify the quality of the generated responses, where win-rates exceeding 50% indicate that the respective preference curation strategy generates responses superior to those of GPT-4V. Additionally, more qualitative case studies are provided in Appendix, offering an intuitive understanding of the quality improvements achieved by TPR. Collectively, all these result confirm that TPR’s meticulous, topic-focused curation not only leads to better performance but also genuinely enhances the intrinsic quality of the preference data for robust VLM alignment.
Rejected Response $y_{l}$. We conduct ablations to assess how different $y_{l}$ curation approaches influence the nature of the resulting $y_{l}$, particularly its fidelity in incorporating hallucinations that reflect the base model’s genuine failure modes. Some rewriting-based methods [27] employ powerful external models, such as GPT-4V, to modify responses, intending to create suitably “negative” $y_{l}$ instances. However, a critical concern is that the hallucination patterns introduced by these external rewriters may deviate from genuine deficiencies of the model that we aim to align. As illustrated in Figure 4 (c), the distribution of hallucination types in $y_{l}$ from these methods can differ significantly from model’s intrinsic hallucination patterns. For instance, an external rewriter might introduce a more imbalanced distribution of various hallucination types, while the model’s own intrinsic hallucination patterns exhibit a relatively uniform distribution. This discrepancy results in a mismatch between the curated $y_{l}$ and model’s actual deficiencies, thereby impairing alignment efficiency (exp 3a vs.3i vs.3j) because the model is not optimally trained against its specific hallucinatory tendencies. Conversely, TPR performs intra-topic replacement on model-generated response template, ensuring that $y_{l}$ more accurately mirror the model’s intrinsic failure distribution.
Data Efficiency. Building on the findings that TPR yields higher quality topic candidates and overall responses, we further investigate its data efficiency, as illustrated in Figure 1 (b). While RLHF-V [13], leveraging human annotation, demonstrates notable initial efficiency by achieving low hallucination rates with only 1.4k data, its prohibitive costs inherently limit scalability. In contrast, TPR, through its automated data curation process, facilitates a rapid reduction in hallucination levels when scaling from 4k to 20k data, ultimately surpassing human-annotated data in both performance and cost-effectiveness. Other automated methods generally exhibit inferior data efficiency. For instance, RLAIF-V [18], despite appearing competitive with 22k data volume, actually necessitates multiple rounds of data re-generation and retraining (effectively equivalent to an 88k data processing effort), significantly inflating its computational cost compared to TPR. These observations underscore how the enhanced data quality, achieved through TPR’s systematic optimization of reward gap configuration, directly contributes to superior data efficiency.
@@
5 Conclusion and Limitations
In this work, we introduced TPR, a novel paradigm designed for robust VLM alignment and visual hallucination mitigation. By systematically controlling the reward gap configuration through topic-level alternatives generation and selective topic rewriting, TPR facilitates the construction of high-quality preference pairs with precisely controlled characteristics. This meticulous data curation is pivotal for generating effective training signals that significantly enhance model performance in mitigating hallucinations. Comprehensive experiments demonstrate that TPR, particularly when augmented with our proposed curriculum learning strategy (TPR-CL), achieves state-of-the-art performance on several hallucination benchmarks, significantly reducing hallucinatory outputs from the reference model.

### v4

@@ 1 Introduction
To address these limitations, we propose Topic-level Preference Rewriting (TPR), a novel VLM hallucination mitigation paradigm designed for systematically optimizing reward gap configuration during data curation, as illustrated in Figure˜1 (a). To provide precise, fine-grained control over semantic details between $y_{w}$ and $y_{l}$, TPR operates at the topic-level. Specifically, TPR first decomposes responses into semantic units and clusters them into distinct topics. It then performs intra-topic self-resampling using the model itself to avoid introducing external biases that may occur with rewriting-based methods. Preference pairs $(y_{w},y_{l})$ are subsequently constructed by selectively replacing original semantic units with these alternatives from the same topic. This mechanism allows for precise, fine-grained adjustment of semantic details between $y_{w}$ and $y_{l}$, guided by various chosen strategies. For example, a greedy strategy can establish highly discriminative reward gaps by constructing $y_{w}$ and $y_{l}$ using the highest- or lowest-scored alternatives from their respective topics.
Moreover, this flexible control over fine-grained differences uniquely enables the formulation and investigation of sophisticated strategies for achieving an optimal reward gap configuration. In this work, we exemplify this capability with a simple yet effective curriculum learning strategy that progressively adjusts the difficulty of hallucinations included in the $y_{l}$. By implementing such hard negative mining, the model can be trained to more effectively counteract subtle and challenging hallucinations, showcasing the effectiveness of TPR. Crucially, this effectiveness is coupled with superior data efficiency, stemming directly from TPR’s ability to curate high-quality preference pairs. This enhanced data efficiency, clearly depicted in Figure˜1 (b), underscores how TPR’s systematic optimization of reward gap configuration leads to more performant and cost-effective VLM alignment.
@@ 3.1 Preliminary: Preference Data and Alignment
Mitigating visual hallucinations in VLMs often involves aligning a policy model $\pi_{\theta}$ with carefully curated preference data, denoted as $\mathcal{D}=\{(I,x,y_{w},y_{l})\}$. Typically, for a given image $I$ and prompt $x$, these methods employ a base reference model $\pi_{\text{ref}}$ to generate candidate responses. Human experts or proxy AI labelers $\pi_{\text{label}}$ then evaluate these responses or rewrite them to form preference pairs $(y_{w},y_{l})$, where $y_{w}$ is preferred over $y_{l}$. The target policy model $\pi_{\theta}$, often initialized from $\pi_{\text{ref}}$, is subsequently fine-tuned on $\mathcal{D}$ using Direct Preference Optimization (DPO) [21].
The core idea of DPO is to optimize the policy $\pi_{\theta}$ to satisfy the preferences in $\mathcal{D}$, which are assumed to follow a latent reward model that reflects the preference probability $p(y_{w}\succ y_{l}|x)$, while simultaneously being constrained by a KL-divergence penalty to not stray too far from the initial reference policy $\pi_{\text{ref}}$. The DPO loss function is formulated as:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
where $\beta$ is a hyperparameter that controls the strength of the preference modeling versus the KL constraint.
As highlighted by recent studies [20, 23, 24], the effectiveness of such alignment critically depends on the true reward gaps, formally defined as the difference $r(y_{w};x)-r(y_{l};x)$ within each preference pair. The strategic configuration of reward gaps across the curated preference data provides substantial training signals that not only reinforce desired behaviors but also accurately expose the model’s genuine and harder-to-discern deficiencies. Therefore, to sculpt effective reward gaps, two key principles should guide data curation: (1) The process must enable fine-grained control over data characteristics in $y_{w}$ and $y_{l}$ to deliberately shape reward gaps. (2) The preference pairs should accurately reflect desired behaviors while exposing model’s intrinsic failure modes.
@@
Guided by these principles, we propose Topic-level Preference Rewriting (TPR), a novel framework designed to systematically configure the reward gap by establishing fine-grained control over topic-level semantic details within the response. TPR implements this through two core steps: topic-level alternatives generation (Section˜3.2) and selective topic replacement (Section˜3.3). Building upon this precise control, TPR enables the active exploration and strategic shaping of reward gap configuration, exemplified by the curriculum learning strategy detailed in Section˜3.4. For clarity, we provide a detailed pseudo code of the complete TPR workflow in Algorithm˜1.
Algorithm 1 Topic-level Preference Rewriting (TPR)
1:Reference model $\pi_{\text{ref}}$, Labeler model $\pi_{\text{label}}$, Source data $\mathcal{D}_{\text{src}}$ (Image II, Prompt $x$), Chosen strategy $\omega$ (e.g., greedy, curriculum).
2:Preference data $\mathcal{D}_{\text{pref}}=\{(I,x,y_{w},y_{l})\}$.
3:Initialize $\mathcal{D}_{\text{pref}}\leftarrow\varnothing$;
4:for each (II, $x$) in $\mathcal{D}_{\text{src}}$ do
5:  Initialize initial responses $S_{y}\leftarrow\varnothing$, semantic units $S_{u}\leftarrow\varnothing$, topic clusters $S_{C}\leftarrow\varnothing$;
6:  for $i\leftarrow 1$ to $M$ do
7:   $y_{i}\leftarrow\text{Sample}(\pi_{\text{ref}},I,x)$;
8:   Add $y_{i}$ to $S_{y}$;
9:   Add $\text{Decompose}(\pi_{\text{ref}},y_{i})$ to $S_{u}$;
10:  end for
11:  $S_{C}\leftarrow\text{TopicCluster}(\pi_{\text{ref}},S_{u})$;
12:  for each cluster $C$ in $S_{C}$ do
13:   $\text{IntraTopicResample}(\pi_{\text{ref}},C)$;
14:   $\text{Rank}(\pi_{\text{label}},C)$;
15:  end for
16:  Initialize response template $y_{k}\leftarrow$ Randomly select from $S_{y}$;
17:  Initialize replacements $S_{w}\leftarrow\varnothing,S_{l}\leftarrow\varnothing$;
18:  for each unit $u_{k}\in C$ in $y_{k}$ do
19:   $(S_{w},S_{l})\leftarrow\text{SelectAlternatives}(\omega,C)$;
20:  end for
21:  $y_{w}\leftarrow\text{InContextRewrite}(\pi_{\text{ref}},y_{k},S_{w})$;
22:  $y_{l}\leftarrow\text{InContextRewrite}(\pi_{\text{ref}},y_{k},S_{l})$;
23:  Add (II, $x$, $y_{w}$, $y_{l}$) to $\mathcal{D}_{\text{pref}}$;
24:end for
25:return $\mathcal{D}_{\text{pref}}$
@@ 3.2 Topic-level Alternatives Generation
VLM responses comprise various semantic topics, encompassing diverse objects and attributes, intricate spatial relationships, or subtle contextual implications. To offer flexible control over fine-grained details within these responses, TPR operates precisely at the topic level. Such a topic-centric approach is made viable by findings like those in RLAIF-V [18], which suggest topics within a response often exhibit weak correlations, permitting their relatively independent manipulation. As illustrated in Figure˜2, TPR introduces a topic-level alternative generation approach designed to provide a rich set of candidates for the subsequent selective replacement of individual semantic topics.
@@
Topic Clustering. To facilitate subsequent selective replacement of semantic units with alternatives under the same topic, we group semantically related units from different candidate responses into unique topic clusters, based on textual consistency and visual correlation. Textual consistency between units $u_{m,n}$ and $u_{p,q}$ is assessed by querying the reference model $\pi_{\text{ref}}$, e.g., “Are $u_{m,n}$ and $u_{p,q}$ describing the same topic?”. For visual correlation, we utilize features from the VLM’s visual encoder (e.g., CLIP [45]) to verify whether both units refer to similar regions within the input image $I$. It is essential for disambiguating textually similar units that describe visually distinct entities. Units $u_{m,n}$ and $u_{p,q}$ are considered the same topic only if they satisfy both textual consistency and visual correlation criteria. We then apply a greedy algorithm [46] for topic clustering, yielding topic cluster $\{c\}$, each containing textually and visually related units. More details are provided in Appendix B.
@@ 3.3 Selective Topic Rewriting
Figure 3: Constructing Preference Pairs by Selectively Replacement. Starting with a pool of scored topic alternatives, preference pairs are constructed through selectively replacing units within the template. The specific alternatives are chosen based on strategies like greedy replacement, to deliberately control over the resulting reward gap. In-context rewriting is then employed to seamlessly integrate these chosen alternatives, ensuring fluent preference pairs for subsequent model alignment.
@@ 3.4 Exploring Optimal Reward Gap Configuration
Leveraging this control, we propose a simple yet effective curriculum learning strategy, analogous to hard negative mining in other domains, to optimize the reward gap configuration for hallucination reduction. This strategy involves structuring the preference data curation and DPO training in stages. Specifically, during an initial “Warm-Up” stage, we adopt the greedy strategy described above, constructing $(y_{w},y_{l})$ pairs with maximal score divergence between corresponding semantic units. This provide a strong initial learning signal, encouraging faster convergence by clearly differentiating between highly discriminative content. In a subsequent “Hard-Mining” stage, we gradually increase the difficulty of the learning objective by constructing $y_{l}$ using incorrect alternative units that have progressively higher scores, i.e., they are less obviously wrong and closer to the decision boundary. Introducing these “hard negative” examples in $y_{l}$ challenges the model to make finer distinctions during DPO training. By learning to differentiate preferred responses from subtly flawed ones, the model can more effectively refine its ability to detect and avoid nuanced hallucinations, ultimately improving robustness. This curriculum strategy systematically shape the overall reward gap configuration over time, starting with broad distinctions and moving towards finer-grained ones ultimately. Further details are provided in Appendix C.
@@ 4.1 Experimental Setup
Models. In line with previous studies, we use LLaVA-1.5-7B [7] as both the reference model $\pi_{\text{ref}}$ for generating preference data, and as the policy model $\pi_{\theta}$ that is subsequently fine-tuned. The labeler model $\pi_{\text{label}}$, used for intra-topic ranking described in Section˜3.3, is LLaVA-NeXT-34B [4], a choice consistent with recent approaches [18, 28]. In our experiments, we evaluate two variants of proposed TPR: (1) TPR utilizes the greedy replacement strategy for data curation. (2) TPR-CL (Curriculum Learning) utilizes a curriculum learning strategy for data curation, detailed in Section˜3.4. We compare TPR against a range of counterparts, including ranking-based methods [16, 18, 19, 20, 25] and rewriting-based methods that rely on human experts [13] or external models [27, 28, 59, 60].
@@
$_{\text{s}} \downarrow CH$$_{\text{i}} \downarrow$ Score $\uparrow$ Hall. $\downarrow$ Acc. $\uparrow$ F1 $\uparrow$ Adv. $\uparrow$ All $\uparrow$ Trust. $\uparrow$ Win. $\uparrow$ Acc. $\uparrow$ Acc. $\uparrow$
@@
Data Source. Following Yu, et al. [13, 18], we curate preference data based on 7 publicly available dataset sources: VQA v2 [47], MSCOCO [48], ShareGPT-4V [49], TextVQA [50], MovieNet [51], OKVQA [52] and Google Landmark v2 [53]. We generate a total of 20,000 preference data instances used for alignment. For the TPR-CL variant, 12,000 instances (60%) are constructed during “Warm-Up” stage, and the remaining 8,000 (40%) are constructed during “Hard-Mining” stage.
@@ 4.2 Main Results
Results in Table˜1 demonstrate TPR significantly mitigates visual hallucinations in VLMs across several benchmarks. The proposed TPR, in both its evaluated variants, achieves leading performance in mitigating visual hallucinations when applied to LLaVA-1.5-7B. Notably, our TPR-CL variant establishes new state-of-the-art results on several hallucination benchmarks, reducing hallucinations by $\sim$93% on Object-HalBench and $\sim$41% on MMHal-Bench. This substantial improvements can be consistently observed across other challenging hallucination benchmarks such as AMBER and RefoMB, underscoring the efficacy of our proposed design. Moreover, models fine-tuned using TPR-curated data not only maintain but sometimes enhance the base model’s performance on general capability benchmarks. This outcome indicates that TPR can effectively suppresses visual hallucinations without compromising general VLM capabilities.
@@
Multi-Res. Decompose Intra-Topic Rsp. Strategy In-Ctx. $_{\text{s}} \downarrow CH$$_{\text{i}} \downarrow$ Acc. $\uparrow$ F1 $\uparrow$
@@
(2d) SR+Both+Greedy
35.5 20.1 80.1 84.4
SR+Both+Greedy 7.2 4.0 80.2 86.4
SR+Both+Greedy 12.6 6.6 77.9 84.9
@@ 4.3 Ablation Studies
To investigate how TPR optimizes the reward gap configuration by enabling deliberate control over underlying data characteristics within preference pairs, we conduct a series of ablation studies. As presented in Table˜2, our investigations focus on addressing the following questions: (1) How do the core components of TPR influence the quality and data characteristics of the resulting preference pairs, which directly shape the individual reward gaps? (2) How does the selective replacement mechanism in TPR impact final performance of the learned policy, when compared against other preference curation approaches, such as ranking-based or rewriting-based methods? (3) How to construct a high-quality rejected response $y_{l}$ in preference pair? Is it more effective to employ external “black-box” rewriters common in rewriting-based methods or to utilize model’s internally resampled candidates, as TPR does? (4) From a broader perspective, what is the data efficiency of preference data curated by TPR, particularly when comparing its performance with that of other methods across varying data volumes? For computational efficiency, these ablation studies are conducted using a subset of 8,000 curated preference data instances. More ablations are provided in the Appendix G, including using different model architectures, using the reference model itself as labeler, and etc.
@@
Topic Alternatives Generation. We conduct ablations on different approaches for generating topic alternatives. Results in Table˜2 reveal that employing multiple initial candidate responses for decomposition and subsequently performing intra-topic self-resampling proves highly beneficial (exp 2a vs.2e/f). It facilitates better coverage of edge cases, such as rarely mentioned attributes, and ultimately yields a more diverse and comprehensive set of alternatives. Such diversity is crucial for sculpting more sophisticated reward gap configuration during data curation. To quantify these aspects, Figure˜4 (b) provides a direct comparison, evaluating the underlying data characteristics like informativeness and trustworthiness of these generated alternatives, based on GPT-4V reviews.
Preference Curation Approaches. We conduct ablations to analyze the impact of specific design choices within the preference curation process, as detailed in Table˜2. Deviations from our base TPR variant (exp 2a), such as employing random selection for replacement (exp 2b), applying selective replacement to only one side of preference pair (exp 2c, or omitting in-context rewriting (exp 2d), all degrade final model performance. These findings highlight that the purposeful, symmetric replacement for both preferred and rejected responses, coupled with in-context rewriting, is crucial for crafting reward gaps that provide strong and effective learning signals for DPO.
Moreover, TPR’s core mechanism of selective topic replacement demonstrably outperforms other preference curation strategies, such as ranking-based or rewriting-based methods (exp 2a vs.2g/h/i/j). These performance advantages are further substantiated by quality assessments present in Figure˜4 (a). In this evaluation, responses generated by different curation strategies are compared against strong “ground truth” (GT) responses from GPT-4V. We calculate one-vs.-one win-rates to quantify the quality of the generated responses, where win-rates exceeding 50% indicate that the respective preference curation strategy generates responses superior to those of GPT-4V. Additionally, more qualitative case studies are provided in Appendix E, offering an intuitive understanding of the quality improvements achieved by TPR. Collectively, all these result confirm that TPR’s meticulous, topic-focused curation not only leads to better performance but also genuinely enhances the intrinsic quality of the preference data for robust VLM alignment.
Rejected Response $y_{l}$. We conduct ablations to assess how different $y_{l}$ curation approaches influence the nature of the resulting $y_{l}$, particularly its fidelity in incorporating hallucinations that reflect the base model’s genuine failure modes. Some rewriting-based methods [27] employ powerful external models, such as GPT-4V, to modify responses, intending to create suitably “negative” $y_{l}$ instances. However, a critical concern is that the hallucination patterns introduced by these external rewriters may deviate from genuine deficiencies of the model that we aim to align. As illustrated in Figure˜4 (c), the distribution of hallucination types in $y_{l}$ from these methods can differ significantly from model’s intrinsic hallucination patterns. For instance, an external rewriter might introduce a more imbalanced distribution of various hallucination types, while the model’s own intrinsic hallucination patterns exhibit a relatively uniform distribution. This discrepancy results in a mismatch between the curated $y_{l}$ and model’s actual deficiencies, thereby impairing alignment efficiency (exp 2a vs.2i vs.2j) because the model is not optimally trained against its specific hallucinatory tendencies. Conversely, TPR performs intra-topic replacement on model-generated response template, ensuring that $y_{l}$ more accurately mirror the model’s intrinsic failure distribution.
Data Efficiency. Building on the findings that TPR yields higher quality topic candidates and overall responses, we further investigate its data efficiency, as illustrated in Figure˜1 (b). While RLHF-V [13], leveraging human annotation, demonstrates notable initial efficiency by achieving low hallucination rates with only 1.4k data, its prohibitive costs inherently limit scalability. In contrast, TPR, through its automated data curation process, facilitates a rapid reduction in hallucination levels when scaling from 4k to 20k data, ultimately surpassing human-annotated data in both performance and cost-effectiveness. Other automated methods generally exhibit inferior data efficiency. For instance, RLAIF-V [18], despite appearing competitive with 22k data volume, actually necessitates multiple rounds of data re-generation and retraining (effectively equivalent to an 88k data processing effort), significantly inflating its computational cost compared to TPR. These observations underscore how the enhanced data quality, achieved through TPR’s systematic optimization of reward gap configuration, directly contributes to superior data efficiency.
5 Limitations and Future Prospects
Our work presents a promising direction for mitigating hallucinations, yet it also has limitations that open avenues for future research.
Complexity of TPR Pipeline. The TPR involves several sequential steps, including response decomposition, topic clustering, self-resampling, and in-context rewriting. While each stage is integral to ensuring high-quality data curation, their combination increases the overall complexity and computational overhead. In future work, we are committed to improving the framework’s efficiency by simplifying the pipeline, e.g., by investigating the necessity of the clustering step, to further enhance its throughput and reduce computational demands.